# FireANTs: Adaptive Riemannian Optimization for Multi-Scale Diffeomorphic Registration

### v1

@@ 2.2 Results on the EMPIRE10 lung CT challenge
We focus on three evaluation critera of the challenge - (1) fissure alignment errors (in %) denoting the fraction of fissure voxels that are misaligned after registration, (2) landmark distance (in mm), and (3) singularity errors - which is defined as the fraction of the image volume that is warped non-diffeomorphically. Results are summarized in  Fig. 3 which also demonstrates the effect of representation choice for modeling diffeomorphisms. For the same scan pairs and cost functions, the DARTEL baseline performs substantially poorly in terms of fissure alignment, landmark distance and singularities than that of ANTs by three orders of magnitude. Our method has about a 5$\times$ lower error than ANTs on the fissure alignment task, and performs better on 5 out of 6 subregions on the landmark distance alignment task. Moreover, although all methods return deformations that are theoretically diffeomorphic, the SVF representation introduces significant signularity errors (voxels where the deformation is not diffeomorphic) due to discretization errors in the Euler integration. The ANTs baseline also introduces some singularities in its proposed diffeomorphic transform. Our method, on the other hand computes numerically perfect diffeomorphic transforms.
@@
Figure 4: Results on the RnR-ExM mouse dataset: (a): As of March 15, 2024, our method ranks first in the mouse brain registration task, which is the only task in the challenge requiring deformable registration. Our top two successful submissions secure the first and second position, with a 0.361 improvement in Dice score compared to the 3rd ranked submission, which is 0.261 better than the 5th ranked submission (bigstream). Note that among the top 10 submissions, our method has the lowest standard deviation (4.42$\times$ lower than the second best submission) showing the robustness of our model across different microscopy volumes. (b): A qualitative comparison of FireANTs with Bigstream  62, the other top leading method in the challenge. The moving image volumes have substantially more noise than the fixed image volumes, making intensity-based registration difficult. The non-rigid deformation dynamics of the hydrogel are clearly visible, as the moving volume has a thicker boundary than the fixed volume. The bigstream baseline does not capture this dynamics very well - the registered volume looks closer to the moving than the fixed volume. Moreover, the affine registration in Bigstream knocks the boundary slices out of the volume, leading to drop in registration performance. On contrary, our method’s affine and deformable stages are more stable, leading to better registration and avoiding spurious out-of-bound artifacts at the boundary slices.
2.3 Evaluation on high-resolution mouse cortex registration
[…]
FireANTs is able to register the volume at native resolution. We perform an affine registration followed by a diffeomorphic registration step. The entire method takes about 2-3 minutes on a single A6000 GPU. As shown in  Fig. 4, our method secures the first place on the leaderboard, with a considerable improvement in the Dice score and a 4.42$\times$ reduction in the standard deviation of the Dice scores compared to the next best method.  Fig. 4 also shows qualitative comparison of our method compared to Bigstream 62, the winner of the RnR-ExM challenge. Bigstream only performs an affine registration, leading to inaccurate registration in one of three test volumes, leading to a lower average Dice score and higher variance. Moreover, the affine registration leads to boundary in-plane slices being knocked out of the volume, leading to poor registration. Our method therefore presents a powerful benchmark for accurate and robust deformable image registration.
@@ 2.4 Ease of experimentation due to efficient implementation
One of the major contributions of our work is to enable fast and scalable image registration while improving accuracy. In applications like atlas/template building, registration is used in an iterative manner (in the ‘inner loop’) of the optimization. Moreover, modalities such as expansion microscopy enable acquisition of high resolution 3D volumes. Another application that requires fast runtimes is hyperparameter tuning, since different datasets and modalities admit notably different hyperparameters for optimal registration. This calls for an increasing need for fast and scalable registration algorithms. To demonstrate the computational and runtime efficiency of our method, we demonstrate the runtime of our library on the brain and lung datasets. All the experiments for our method are run on a single A6000 GPU, with a batch size of 1 (to avoid amortizing the time over a bigger batch size). For the brain datasets, we run ANTs with the recommended configuration with AMD EPYC 7713 Processor (single thread) and 512GB RAM. For the EMPIRE10 lung dataset, we use the runtimes described in the writeup provided as part of the challenge. A runtime analysis of our method on the brain and EMPIRE10 datasets are shown in Fig. 5.
(a) Timing analysis on four brain MRI datasets.
@@
(b) Timing analysis on EMPIRE10 dataset
@@
Figure 5: Timing analysis of our library: We compare the runtime of our implementation with the ANTs library. (a) shows distribution of speedup (runtime of ANTs divided by runtime of our method) and statistics of runtimes (in seconds) for the four brain MRI datasets. For all datasets, our implementation runs a minimum of two orders of magnitudes faster, making it suitable for hyperparameter search algorithms, and larger datasets. Table (b) shows the runtime of ANTs, DARTEL and our implementation on the EMPIRE10 challenge data. The first three colums show the actual runtime of the methods, followed by the speedup obtained by our method when compared to ANTs and DARTEL. Note that our method runs a minimum of 320 times faster than ANTs, saving a substantial amount of time, at no loss in registration quality.
@@
Figure 6: Feasibility of hyperparameter searches on LPBA40 and EMPIRE10 datasets: The speed of FireANTs makes hyperparameter studies like these feasible, which ANTs would take years to complete. (a): We perform a hyperparameter grid search on three hyperparameters of interest - smoothing kernel for the warp field ($\sigma_{\text{warp}}$) in pixels, smoothing kernel for the gradient of warp field ($\sigma_{\text{grad}}$) in pixels and learning rate $\eta$. The metric to optimize in this case is the target overlap. For the LPBA40 dataset, we perform a hyperparameter sweep over 640 configurations in 40 hours with 8 A6000 GPUs. A corresponding hyperparameter sweep with 8 concurrent jobs with each job consuming 8 CPUs would take $\sim$3.6 years to complete. The white contour representing the level set for target overlap = 0.75, and the black contour representing the level set for target overlap of 0.74 show the robustness of our method to hyperparameters - performance is not brittle or sensitive to choice of hyperparameters. (b): Hyperparameter grid search on the EMPIRE10 dataset over $\sigma_{\text{warp}}$ and $\sigma_{\text{grad}}$ parameters (456 configurations), with a fixed learning rate of $\eta=0.25$. The metric to optimize is the Dice score of the provided binary lung mask. This sweep takes about 12.37 hours on 8 GPUs, whereas a corresponding sweep would take 296 days for ANTs and 345 days for DARTEL (more in Fig. 5). The white contour corresponds to the level set for Dice score = 0.96, showing both a huge spectrum of parameters that achieve high Dice scores, and low sensitivity of the method to choice of hyperparameters.
@@
For the EMPIRE10 dataset, our method reduces the runtime from 1 to 12 hours for a single scan pair to under a minute. We compare our method with both ANTs and DARTEL implementations. Since the exponential map requires a few integration steps for each iteration, this variant is even slower than ANTs. Our method enjoys a minimum of more than 300$\times$ speedup over ANTs. On the brain datasets, our method achieves a consistent speedup of 3 orders of magnitude. This happens due to a better choice of hyperparameters compared to the baseline, faster convergence due to the adaptive optimization, and better memory and compute utilization by cuDNN implementations. These improvements in runtime occur while also providing at par, or superior results (Fig. 23, Extended Data S.4S.5).
2.4.1 Fast hyperparameter tuning using FireANTs
@@
In optimization toolkits such as ANTs, several hyperparameters are key to high quality registration. Some of these hyperparameters are the window size for the similarity metric Cross-Correlation or bin size for Mutual Information. In our experience, the Gaussian smoothing kernel $\sigma_{\text{grad}}$ for the gradient of the warp field are two of the most important parameters for diffeomorphic registration. The optimal values of these hyperparameters vary by image modality, intensity profile, noise and resolution. Typically, these values are provided by some combination of expertise of domain experts and trial-and-error. However, non experts may not be able to adopt these parameters in different domains or novel acquisition settings. Recently, techniques such as hyperparameter tuning have become popular, especially in deep learning. In the case of registration, hyperparameter search can be performed by considering some form of label/landmark overlap measure between images in a validation set. We demonstrate the stability and runtime efficiency of our method using two experiments : (1) Owing to the fast runtimes of our implementation, we show that hyperparameter tuning is now feasible for different datasets. The optimal set of hyperparameters is dependent on the dataset and image statistics, as shown in the LPBA40 and EMPIRE10 datasets; (2) within a particular dataset, the sensitivity of our method around the optimal hyperparameters is very low, demonstrating the robustness and reliability of our method.
@@
We choose the LPBA40 dataset among the 4 brain datasets due to its larger size (40$\times$39 = 1560 pairs). We choose three parameters to search over : the learning rate ($\eta$), and the gaussian smoothing parameters $\sigma_{\text{warp}},\sigma_{\text{grad}}$. We use the Ray library (https://docs.ray.io/) to perform a hyperparameter tuning using grid search. For the LPBA40 dataset, a grid search over three parameters (shown in  Fig. 6) takes about 40.4 hours with 8 parallel jobs. On the contrary, ANTs would require around 3.6 years to complete the same grid search, with 8 threads allocated to each job and 8 parallel jobs. This makes hyperparameter search for a unknown modality computationally feasible.  6(a) shows a dense red region suggesting the final target overlap is not sensitive to the choice of hyperparameters. Specifically, the maximum target overlap is 0.7565 and 58.4% of these configurations have an average target overlap of $\geq$ 0.74. This is demonstrated in  Fig. 6 (top) by the white contour line denoting the level set for target overlap = 0.75, and the black contour line denoting the level set for target overlap of 0.74. The target overlap is quite insensitive to the learning rate ($\geq 0.4$) showing that our algorithm achieves fast convergence with a smaller learning rate. On the EMPIRE10 dataset, we fix the learning rate and perform a similar hyperparamter search over two parameters, the Gaussian smoothing parameters $\sigma_{\text{warp}},\sigma_{\text{grad}}$. We use the average Dice score between the fixed and moving lung mask to choose the optimal hyperparameters. FireANTs can perform a full grid search over 456 configurations on the EMPIRE10 dataset in 12.37 hours with 8 A6000 GPUs, while it takes SyN 10.031 days to run over a single configuration, or equivalently around 296 days for the entire grid search. Normalizing for 8 concurrent jobs and 456 configurations, it would take ANTs about 296 days, and DARTEL about 345 days. This shows that our method and accompanying implementation can now make hyperparameter search for 3D image registration studies feasible.
@@ 3 Discussion
We present FireANTs, a powerful and general purpose multi-scale registration algorithm. Our method performs registration by generalizing the concept of first-order adaptive optimization schemes for optimizing parameters in the Euclidean space, to diffeomorphisms. This is highly non-trivial because diffeomorphisms are typically implemented as an image grid proportional to the size of the fixed image, and are optimized in a multi-scale manner to capture large deformations 32, 60, 53 leading to changing grid size over the course of optimization. Our method also avoids a computationally expensive parallel transport step for diffeomorphisms by solving a modified instance of registration at each time step. Our method achieves consistent improvements in performance over state-of-the-art optimization-based registration algorithms like ANTs, DARTEL, SynthMorph and Bigstream. This improvement is shown across six datasets with a spectrum of anatomies (in-vivo human brain, human and ovine lungs, mouse cortex), contrast, image volume sizes (ranging from 196 up to 2048 voxels per dimension), and modalities (MRI, CT, microscopy). A key advantage of our method is that we do not tradeoff any of accuracy, speed or robustness for the others, thus being a powerful registration algorithm.
@@
Our method shows consistent improvements and robust performance on four public brain MRI datasets. Many classical image registration methods have been developed for neuroimaging studies 60, 53, 38, 70 and registration still remains a crucial component in brain mapping 71, 72. FireANTs’ consistent improvement in performance can be attributed to the quasi-second-order update which normalizes the varying curvature of the per-pixel gradient, leading to faster convergence. This performance is consistent across metrics ( Fig. 2) and anatomical structures ( Fig. S.1). With acquisition of larger datasets 73 and high-resolution imaging 74, fast and accurate registration runtimes become imperative to enable large-scale studies. Our performance comes with a reduction of runtime of up to three orders of magnitude.
@@
We also demonstrate competitive performance in the EMPIRE10 challenge, widely regarded as a comprehensive evaluation of registration algorithms 63, 75. Unlike the brain imaging datasets, the EMPIRE10 dataset contains images with large deformations, anisotropic image spacings and sizes, and thin structures like airways and pulmonary fissures which are hard to align based on image intensity alone. These image volumes are typically much larger than what deep learning methods can currently handle at native resolution 75, 52, 46. FireANTs performs much better registration in terms of landmark, fissure alignment and singularities, while being two orders of magnitude faster. This experiment also calls attention to a much overlooked detail - the performance gap due to choice of representation of diffeomorphisms (direct Riemannian optimization versus exponential map). We show that direct Riemannian optimization is preferable to exponential maps, both in FireANTs and in baselines (ANTs versus DARTEL). This is because computing the exponential map is expensive for diffeomorphisms, the number of iterations can be large for large deformations 53. For example, in Fig. 5(a), the exponential map representation (DARTEL) takes substantially longer to run, compared to ANTS. Shooting methods modify the velocity field at each iteration and tend to be sensitive to hyper-parameter choices. For example, in Fig. 3 the results for shooting methods are substantially worse those for methods that optimize the transformation directly. We also observe this for the LPBA40 dataset in Fig. S.6, where over a wide range of hyperparameter choices, the shooting method consistently under-performed.
Our method is consistently 300–2000$\times$ faster, is robust to choice of hyperparameters, allowing users to utilize principled hyperparameter search algorithms for novel applications or modalities. This opens up many other avenues in registration, including modalities such as microscopy, ex-vivo imaging where advanced imaging techniques have led to ultra high resolution data acquisition protocols. Registration algorithms are crucial to subsequent downstream tasks in these studies, it is therefore imperative for registration algorithms to be accurate and scale with the data as well. Our method showcases this on the RnR-ExM mouse cortex dataset, where our method performs the best overall in a 2-3 minute runtime on a single GPU. Our method and accompanying implementation is a step towards this avenue, making advanced registration algorithms fast and accessible.
@@
In summary, FireANTs is a powerful and general purpose multi-scale registration algorithm and sets a new state-of-the-art benchmark. We propose to leverge the accurate, robust and fast library to speed up registration workflows for modalities like microscopy and ex-vivo imaging (humans, mice, C. Elegans, etc.), where imaging resolutions are large and algorithms are bottlenecked by scalability.

### v2

@@ 2.2 Results on state-of-the-art biomedical benchmarks
We consider three evaluation criteria of the challenge - (1) fissure alignment errors (in %) denoting the fraction of fissure voxels that are misaligned after registration (Figs. 2 and 2), (2) landmark distance (in mm) (Fig. 2), and (3) singularity errors - which is defined as the fraction of the image volume that is warped non-diffeomorphically (Fig. 2). Results in  Fig. 2 also demonstrates the effect of representation choice for modeling diffeomorphisms. For the same scan pairs and cost functions, the exponential map-based DARTEL performs substantially poorly in terms of fissure alignment, landmark distance and singularities than that of ANTs by three orders of magnitude. In contrast, our method has about a 5$\times$ lower error than ANTs on the fissure alignment task, and performs better on 5 out of 6 subregions on the landmark distance alignment task. Moreover, although all methods return deformations that are theoretically diffeomorphic, the SVF representation introduces significant singularity errors (voxels where the deformation is not diffeomorphic) due to non-adaptive value of $N$ in the scaling-and-squaring step. The ANTs baseline also introduces some singularities in its proposed diffeomorphic transform. Our method, on the other hand computes numerically perfect diffeomorphic transforms. Finally, we compare the fissure alignment error of all submissions in the EMPIRE10 challenge, and show the top 10 algorithms in  Fig. 2. Results demonstrate that FireANTs attains the lowest landmark alignment errors compared to an array of contemporary state-of-the-art algorithms. FireANTs also has the lowest runtime of all 10 methods, signifying a paradigm shift in the balance between computational efficiency and accuracy instigated by our approach.
2.3 Evaluation on high-resolution mouse cortex registration
[…]
In contrast, FireANTs is able to register the volume at native resolution. We perform an affine registration followed by a diffeomorphic registration step. The entire method takes about 2-3 minutes on a single A6000 GPU. As shown in  Fig. 3, our method secures the first place on the leaderboard, with a considerable improvement in the Dice score and a 4.42$\times$ reduction in the standard deviation of the Dice scores compared to the next best method.  Fig. 3 also shows qualitative comparison of our method compared to Bigstream 38, the winner of the RnR-ExM challenge. Bigstream only performs an affine registration, leading to inaccurate registration in one of three test volumes, leading to a lower average Dice score and higher variance. Moreover, the affine registration leads to boundary in-plane slices being knocked out of the volume, leading to poor registration (Fig. 3). FireANTs preserves the boundary in-plane slices during its affine step, and subsequently performs an accurate diffeomorphic registration. This shows the versatility and applicability of FireANTs for high-resolution microscopy registration.
@@ 2.3 Evaluation on high-resolution mouse cortex registration
Figure 3: FireANTs secures first rank in the RnR-ExM mouse dataset: (a): As of January 17, 2025, our method ranks first in the mouse brain registration task, which is the only task in the challenge requiring deformable registration. Our top two successful submissions secure the first and second position, with a 0.361 improvement in Dice score compared to the 3rd ranked submission, which is 0.261 better than the 5th ranked submission (bigstream). Note that among the top 10 submissions, our method has the lowest standard deviation (4.42$\times$ lower than the second best submission) showing the robustness of our model across different microscopy volumes. (b) shows a qualitative comparison of FireANTs with Bigstream  38, the other top leading method in the challenge. The moving image volumes have substantially more noise than the fixed image volumes, making intensity-based registration difficult. The non-rigid deformation dynamics of the hydrogel are clearly visible, as the moving volume has a thicker boundary than the fixed volume. Bigstream does not capture these dynamics very well – this is illustrated by comparing the thickness of the cortex at various points (zoomed orange crops in bottom row), where Bigstream does not deform the cortex enough to match the fixed image. FireANTs deforms and accurately depicts these morphological changes, which can be crucial for downstream morphometric studies. Moreover, the affine registration in Bigstream knocks the boundary slices out of the volume (red highlight in top row), leading to drop in registration performance. On contrary, our method’s affine and deformable stages are more stable, leading to better registration and avoiding spurious out-of-bound artifacts at the boundary slices.
@@ 2.4 Ease of experimentation due to efficient implementation
One of the major contributions of our work is to enable fast and scalable image registration while improving accuracy. In applications like atlas/template building, registration is used in an iterative manner (in the ‘inner loop’) of the optimization. Another application that requires fast runtimes is hyperparameter tuning, since different datasets and modalities admit notably different hyperparameters for optimal registration. This calls for an rapidly escalating need for fast and scalable registration algorithms. To demonstrate the computational and runtime efficiency of our method, we demonstrate the runtime of our library on the brain and lung benchmarks. All the experiments for our method are run on a single A6000 GPU, with a batch size of 1, to avoid amortizing runtime over a bigger batch size. In practice, higher speedups can be obtained by registering batches of images that fit in a single GPU. For the brain datasets, we run ANTs with the recommended configuration with AMD EPYC 7713 Processor (single thread) and 512GB RAM. For the EMPIRE10 lung dataset, we use the runtimes described in the writeup provided as part of the challenge. A runtime analysis of our method on the brain and EMPIRE10 datasets are shown in Fig. 4.
(a) Histogram and summary of runtime on four brain MRI datasets.
@@
(b) Runtime analysis and summary on EMPIRE10 dataset
@@
Figure 4: FireANTs runs with unprecedented speedups. We compare the runtime of our implementation with the ANTs library. (a) shows histogram of speedup (runtime of ANTs divided by runtime of our method) and statistics of runtimes (in seconds) for the four brain MRI datasets. For all datasets, our implementation runs a minimum of two orders of magnitudes faster, making it suitable for hyperparameter search algorithms, and larger datasets. Table (b) shows the runtime of ANTs, DARTEL and our implementation on the EMPIRE10 challenge data. The first three colums show the actual runtime of the methods, followed by the speedup obtained by our method when compared to ANTs and DARTEL. Note that our method runs a minimum of 320 times faster than ANTs, saving a substantial amount of time, at no loss in registration quality.
@@
Figure 5: FireANTs facilitates feasibilility of extensive hyperparameter search in registration The speed of FireANTs makes hyperparameter studies like these feasible, which ANTs would take years to complete. (a): We perform a hyperparameter grid search on three hyperparameters of interest - smoothing kernel for the warp field ($\sigma_{\text{warp}}$) in pixels, smoothing kernel for the gradient of warp field ($\sigma_{\text{grad}}$) in pixels and learning rate $\eta$. The metric to optimize in this case is the target overlap. For the LPBA40 dataset, we perform a hyperparameter sweep over 640 configurations in 40 hours with 8 A6000 GPUs. A corresponding hyperparameter sweep with 8 concurrent jobs with each job consuming 8 CPUs would take $\sim$3.6 years to complete. The white contour representing the level set for target overlap = 0.75, and the black contour representing the level set for target overlap of 0.74 show the robustness of our method to hyperparameters - performance is not brittle or sensitive to choice of hyperparameters. (b): Hyperparameter grid search on the EMPIRE10 dataset over $\sigma_{\text{warp}}$ and $\sigma_{\text{grad}}$ parameters (456 configurations), with a fixed learning rate of $\eta=0.25$. The metric to optimize is the Dice score of the provided binary lung mask. This sweep takes about 12.37 hours on 8 GPUs, whereas a corresponding sweep would take 296 days for ANTs and 345 days for DARTEL (more in Fig. 4). The white contour corresponds to the level set for Dice score = 0.96, showing both a huge spectrum of parameters that achieve high Dice scores, and low sensitivity of the method to choice of hyperparameters.
FireANTs shows up to three orders faster runtime than ANTs
@@
For the EMPIRE10 dataset, we compare our method with both ANTs and DARTEL implementations. We capture the runtime for each scan pair and aggregate the average, minimum and maximum runtimes and speedups in Fig. 4. FireANTs reduces the runtime from 1 to 12 hours for a single scan pair to under a minute. Since the exponential map requires a few integration steps for each iteration and pullback of the gradient from the warp to the velocity field, this variant is significantly slower than ANTs. In contrast, our method achieves a minimum of more than 300$\times$ speedup over ANTs, while maintaining high accuracy. In the brain datasets, our method achieves a consistent speedup of three orders of magnitude. This occurs due to a better choice of hyperparameters compared to the baseline, faster convergence due to the adaptive optimization, and better memory and compute utilization by efficient cuDNN implementations.
On the brain datasets, we compare our method with ANTs, and compute the runtime and speedups for each scan pair. We then plot a histogram of the speedup and aggregate the runtime and speedup in Fig. 4. On all datasets, the average speedup is over 1000$\times$, signifying a new paradigm for image registration workflows. These improvements in runtime occur while also providing at par, or superior quantitative and qualitative results (Fig. 12, Extended Data 1314).
FireANTs allows fast hyperparameter tuning
@@
In optimization toolkits such as ANTs, correct choice of hyperparameters are key to high quality registration. Some of these hyperparameters are the window size for the similarity metric Cross-Correlation or bin size for Mutual Information. In our experience, the Gaussian smoothing kernel $\sigma_{\text{grad}},\sigma_{\text{warp}}$ for the gradient and the warp field are two of the most important parameters for accurate diffeomorphic registration. The optimal values of these hyperparameters vary by image modality, intensity profile, noise and resolution. Typically, these values are provided by some combination of expertise of domain experts and trial-and-error. However, non experts may not be able to adopt these parameters in different domains or novel acquisition settings. Recently, techniques such as hyperparameter tuning have become popular, especially in deep learning.
@@
In the case of registration, hyperparameter search can be performed by considering some form of label/landmark overlap measure between images in a validation set. We demonstrate the stability and runtime efficiency of our method using two experiments : (1) Owing to the fast runtimes of our implementation, we show that hyperparameter tuning is now feasible for different datasets. The optimal set of hyperparameters is dependent on the dataset and image statistics, as shown in the LPBA40 and EMPIRE10 datasets; (2) within a particular dataset, the sensitivity of our method around the optimal hyperparameters is very low, demonstrating the robustness and reliability of our method. We choose the LPBA40 dataset among the 4 brain datasets due to its larger size (40$\times$39 = 1560 pairs). We choose three parameters to search over : the learning rate ($\eta$), and the gaussian smoothing parameters $\sigma_{\text{warp}},\sigma_{\text{grad}}$. We use the Ray library (https://docs.ray.io/) to perform a hyperparameter tuning using grid search. For the LPBA40 dataset, a grid search over three parameters (shown in  Fig. 5) takes about 40.4 hours with 8 parallel jobs. On the contrary, ANTs would require around 3.6 years to complete the same grid search, with 8 threads allocated to each job and 8 parallel jobs. This makes hyperparameter search for an unknown modality computationally tractable. Fig. 5 shows a dense red region suggesting the final target overlap is not sensitive to the choice of hyperparameters. Specifically, the maximum target overlap is 0.7565 and 58.4% of these configurations have an average target overlap of $\geq$ 0.74. This is demonstrated in Fig. 5 by the white contour line denoting the level set for target overlap = 0.75, and the black contour line denoting the level set for target overlap of 0.74. The target overlap is quite insensitive to the learning rate ($\geq 0.4$) showing that our algorithm achieves fast convergence with a smaller learning rate. On the EMPIRE10 dataset, we fix the learning rate and perform a similar hyperparameter search over two parameters, the Gaussian smoothing parameters $\sigma_{\text{warp}},\sigma_{\text{grad}}$, shown in Fig. 5. We use the average Dice score between the fixed and moving lung mask to choose the optimal hyperparameters. FireANTs can perform a full grid search over 456 configurations on the EMPIRE10 dataset in 12.37 hours with 8 A6000 GPUs, while it takes SyN 10.031 days to run over a single configuration. Normalizing for 8 concurrent jobs and 456 configurations, it would take ANTs about 296 days, and DARTEL about 345 days. This shows that our method and accompanying implementation can now make hyperparameter search for 3D image registration studies feasible.
@@ 3 Discussion
We present FireANTs, a powerful and general-purpose multi-scale registration algorithm. Our method performs registration by generalizing the concept of first-order adaptive optimization schemes for optimizing parameters in a fixed Euclidean space, to multi-scale diffeomorphisms. This generalization is highly non-trivial because diffeomorphisms are typically implemented as an image grid proportional to the size of the fixed image, and are optimized in a multi-scale manner to capture large deformations 40, 13, 17 leading to changing grid size throughout optimization. Our method also avoids computationally expensive parallel transport and riemannian metric tensor computation steps for diffeomorphisms by solving an Eulerian descent that exploits the group structure to define descent directions from the identity transform. FireANTs achieves consistent improvements in performance over state-of-the-art optimization-based registration algorithms like ANTs, DARTEL, SynthMorph and Bigstream. This improvement is shown across six datasets with a spectrum of anatomies (in-vivo human brain, human, ovine and synthetic lungs, mouse cortex), contrast, image volume sizes (ranging from 196 up to 2048 voxels per dimension), and modalities (MRI, CT, microscopy). A key advantage of our method is that we do not tradeoff any of accuracy, speed, or robustness for the others, thus being a powerful registration algorithm.
@@
Our method shows consistent improvements and robust performance on four community reference brain MRI datasets. Researchers have developed numerous classical image registration methods specifically for neuroimaging studies 13, 17, 41, 42 but registration remains an open challenge in brain mapping 43, 44. FireANTs’ consistent improvement in performance can be attributed to the quasi-second-order update, which normalizes the varying curvature of per-pixel gradients, leading to faster convergence and better local minima. This performance is consistent across metrics (Fig. 1) and anatomical structures (Fig. 10). With the acquisition of larger datasets 45 and high-resolution imaging 27, 46, 32 owing to rapidly growing image acquisition technology and innovations, fast runtime and accurate registration become imperative to enable large-scale studies. Our performance comes with a reduction of runtime of up to three orders of magnitude on standard datasets.
@@
We also demonstrate competitive performance in the EMPIRE10 challenge, widely regarded as a comprehensive evaluation of registration algorithms 47, 48 for lung CT images. Unlike the brain imaging datasets, the EMPIRE10 dataset contains images with large deformations, anisotropic image spacings and voxel sizes, thin structures like airways and pulmonary fissures, which are hard to align based on image intensity alone. These image volumes are typically much larger than what deep learning methods can currently handle at native resolution 48, 39, 30. FireANTs performs much better registration in terms of landmark, fissure alignment, and singularities, while being two orders of magnitude faster than existing state-of-the-art. This experiment also calls attention to a much-overlooked detail - the performance gap due to the choice of representation of diffeomorphisms (direct optimization versus exponential map). We show that direct optimization of the group element is preferable to exponential maps, both in FireANTs and in baselines (ANTs versus DARTEL). This improvement can be attributed to the representation - one can interpret direct optimization as integrating a set of time-dependent velocity fields since the gradients change throughout optimization, allowing more flexibility in the space of diffeomorphisms it can represent, whereas SVF performs the integral of a time-independent velocity field by design. Moreover, computing the exponential map is expensive for diffeomorphisms since the number of iterations can be higher for larger deformations 17. For example, in Fig. 4(a), the exponential map representation (DARTEL) takes substantially longer to run than ANTs. Shooting methods modify the velocity field at each iteration and tend to be sensitive to hyper-parameter choices. For example, in Fig. 2, the results for shooting methods are substantially worse than for methods that optimize the transformation directly. We also observe this for the LPBA40 dataset in Fig. 15, where the shooting method consistently underperformed over a wide range of hyperparameter choices.
FireANTs is consistently 300–2000$\times$ faster than existing state-of-the-art, is robust to the choice of hyperparameters, allowing users to utilize principled hyperparameter search algorithms for novel applications or modalities. This paradigm shift in runtime will enable fast and accurate registration of high-resolution mesoscale and microscale imaging data that will play a paramount role in advancing our understanding of connectomics, neuroscience, cellular and molecular biology, genetics, pathology, among many other disciplines in the biomedical and biological sciences. With breakthrough advances in high-resolution, high-throughput imaging techniques, it is imperative for registration algorithms to also scale with the inordinate amounts of data.
@@
In summary, FireANTs is a powerful and general-purpose multi-scale registration algorithm and sets a new state-of-the-art benchmark. We propose to leverage the accurate, robust, and fast library to speed up registration workflows for ever-growing needs of performant and fast image registration in a spectrum of disciplines within the biomedical and biological sciences, wherein algorithms are bottlenecked by scalability.